Circular No. 6174

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995N IN MCG –02-38-017 P. Garnavich and P. Challis, Center for Astrophysics, write: “A spec-

trogram of SN1995N was obtained by P. Berlind with the 1.5-m Tillinghast telescope on May 24.3 UT. In addition to the emission lines listed by Benetti et al. (IAUC 6170), we identify unresolved lines of [Ne III] at 389.2 and 399.3 nm; [O III] at 439.1 nm; He II at 471.6 nm; and [O III] at 499.1 and 504.0 nm. These last two [O III] lines show a strong, broad component with a halfwidth of 4500 km/s. However, no corresponding broad emission is detected around [O III] at 439.1 nm. The flux ratio of the auroral and nebular [O III] lines implies a density in the narrow line region of > 10^{6} electrons cm^{−3}.”

NOVA AQUILAE 1995 R. W. Argyle and L. V. Morrison, Royal Greenwich Observatory, report

an accurate optical position for N Aql 1995, obtained with the Carlsberg Automatic Meridian Circle on La Palma and operated by O. Einicke, C. Fabricius, and R. Villamil. The position for equinox J2000.0, FK5 reference frame (epoch 1995.37) is α = 19^{h}05^{m}26^{s}.636 ± 0^{s}.004, δ = −1^{o}42′03′′.27 ± 0′′.07, from 10 observations.

X-RAY NOVA 1993 IN OPHIUCHUS E. A. Karitskaya, Institute of Astronomy, Russian Academy of Sci-

ences; and V. P. Goranskij, Sternberg State Astronomical Institute, Moscow, communicate: “V2293 Oph = GRS 1716–249 is optically bright again. Image-tube photometry with the 1-m telescope of the Tien-Shan High-Altitude Observatory, Kazakhstan, yields: Apr. 24.875 UT, V = 17.6; 30.893, 17.7; May 10.917, 17.2 (accuracy ±0.3).”

TU LEONIS = (8) FLORA L. Schmadel, Astronomisches Rechen-Institut, Heidelberg; P. Schmeer,

Bischmisheim; and F. Börngen, Thüringer Landessternwarte, Tautenburg, write that the presumed variability of TU Leo (α = 9^{h}27^{m}.0, δ = +21^{o}37′, equinox 1950.0; suspected U-Gem type) seems to be entirely attributed to the single observation of an apparent brightening to B = 11 from its normal state as a G-type star of B = 15. They note that this observation, made at the Moscow Observatory on 1917 Mar. 25.8 UT, was evidently of the minor planet (8) Flora, which would then have been within 1′ of the star.

1995 May 26 Daniel W. E. Green